The energy densities of various types of coal are listed below. Anthracite 35 kJ/g Subbituminous 31 kJ/g



Bituminous 28 kJ/g Lignite 26 kJ/g



An unknown sample of one of these coals is burned in an apparatus with a calorimeter constant of 1.3 kJ/°C. When a



0.367-g sample is used, the temperature change is 8.75°C. Which type of coal is the sample?

1
Expert's answer
2021-11-11T14:46:02-0500

Calculate the heat generated by the calorimeter.

"q_{cal} = c_{cal}ΔT \\

= 1.3 \times 8.75 \\

= 11.375 \; kJ"

Calculate the heat released by the coal.

"q_{coal} = -q_{cal} \\

= -11.375 \;kJ"

Calculate the energy density.

"ED = \frac{-q_{coal}}{m_{coal}} \\

= \frac{-(-11.375 \; kJ)}{0.375 \;g} \\

= 31 \; kJ/g"

Therefore, we conclude that the type of coal is subbituminous.